OpenStudy (anonymous):

Use L'hospital rule to find the following limit: Lim x-->0+ 6(tan(4x))^x I did this: Lim x-->0+ xln 6(tan(4x)) which is the same as: ln 6(tan(4x)) / (1/x) Take the derivative: (1/6tan(4x))) * 6(sec^2(4x))(4) / (-1/x^2). But I can't find a way to cancel out my denominator :/ the answer isn't infinity either

OpenStudy (anonymous):

leave the six out of it, put it in at the end

OpenStudy (anonymous):

\[x\ln(\tan(4x)\] \[\frac{\ln(\tan(4x))}{\frac{1}{x}}\]etc

OpenStudy (anonymous):

\[-4x^2\csc(4x)\sec(4x)\] i think is the next step

OpenStudy (anonymous):

oh i see still a problem

OpenStudy (anonymous):

maybe flip the other way

OpenStudy (sirm3d):

\[\large= \frac{ 4x }{ \sin 4x }\frac{ -x }{ \cos 4x }\] there is no problem

OpenStudy (anonymous):

your sin will make the denominator 0 even if multiplied :/

OpenStudy (anonymous):

or am I just not understanding something ?

OpenStudy (anonymous):

no @sirm3d as it! first limit is 1

OpenStudy (anonymous):

*has it

OpenStudy (anonymous):

0ooh take the limit of each... stupid me.

OpenStudy (anonymous):

Thanks to both of you